Calc+ Company manufactures calculators for schools. The master budget is based on sales of 40,000 units at $65 per calculator. Budgeted variable costs are $45 per unit, while budgeted fixed costs total $670,000. Actual income was $211,000 on actual sales of 42,000 units at $64 each. Actual variable costs were $43 per unit and actual fixed costs totaled $671,000. What is the master-budget variance of operating income (list variance amount and if it is favorable or unfavorable)?

Calc+ Company manufactures calculators for schools. The master budget is based on sales of 40,000 units at $65 per calculator. Budgeted variable costs are $45 per unit, while budgeted fixed costs total $670,000. Actual income was $211,000 on actual sales of 42,000 units at $64 each. Actual variable costs were $43 per unit and actual fixed costs totaled $671,000. The flexible budget will report ________ for operating income.

Calc+ Company manufactures calculators for schools. The master budget is based on sales of 40,000 units at $65 per calculator. Budgeted variable costs are $45 per unit, while budgeted fixed costs total $670,000. Actual income was $211,000 on actual sales of 42,000 units at $64 each. Actual variable costs were $43 per unit and actual fixed costs totaled $671,000. What is the spending variance of sales (list variance amount and if it is favorable or unfavorable)?

Calc+ Company manufactures calculators for schools. The master budget is based on sales of 40,000 units at $65 per calculator. Budgeted variable costs are $45 per unit, while budgeted fixed costs total $670,000. Actual income was $211,000 on actual sales of 42,000 units at $64 each. Actual variable costs were $43 per unit and actual fixed costs totaled $671,000. What is the volume variance of variable costs (list variance amount and if it is favorable or unfavorable)?
Answer #1
Master budget variance of operating income:
Master
Budget
Actual
Based on
40000
units
Based on
42000
units
Sales 2600000 2688000
(40000*65) (42000*64)
Less: Variable cost 1800000 1806000
(40000*45) (42000*43)
Contribution margin 800000 882000
Less: Fixed cost 670000 671000
Operating income 130000 211000
Actual operating income is more than the budgeted operating income.Hence, variance is favorable
Master-budget variance of operating income=211000-130000=$81000 favorable
Flexible budget is prepared based on actual units sold:
Actual
Based on
42000
units
Sales 2730000
(42000*65)
Less: Variable cost 1890000
(42000*45)
Contribution margin 840000
Less: Fixed cost 670000
Operating income 170000
The flexible budget will report $ 170000 for operating income.
Spending variance of sales=Actual units sold*(Actual selling price-Budgeted selling price)
If the answer is positive,variance is favorable.Otherwise unfavorable
Spending variance of sales=42000*(64-65)=42000*(-1)=-42000=42000 unfavorable
Volume variance of variable cost=Budgeted variable cost rate*(budgeted units-Actual units )
If the answer is positive,variance is favorable.Otherwise unfavorable
Volume variance of variable cost=45*(40000-42000)=45*(-2000)=-90000=$ 90000 unfavorable